# VasoMIM: Vascular Anatomy-Aware Masked Image Modeling for Vessel Segmentation

### v1

@@ Conclusion
Pre-Training. (1) ARCADE (Popov et al. 2024): This dataset comprises two subsets for coronary vessel segmentation and atherosclerotic plaque localization. Each subset contains $1,500$ images, split into $1,000$ for training, $200$ for validation and $300$ for testing. Only the $1,000$ training images per subset are used for pre-training. (2) CADICA (Jiménez-Partinen et al. 2024): It comprises coronary X-ray angiography videos from $42$ patients, with frames of $512\times 512$ pixels and video lengths ranging from $1$ to $151$ frames. We select $6,594$ high-quality frames from these videos for pre-training. (3) Stenosis (Danilov et al. 2021): It offers X-ray angiography videos from $100$ patients with a total of $8,325$ frames. We use the $7,492$ training frames (from $70$ patients) for pre-training. (4) SYNTAX (Mahmoudi et al. 2025): This dataset includes $2,943$ X-ray angiograms from $231$ patients. All images are used for pre-training. (5) XCAD (Ma et al. 2021): This dataset provides $1,747$ X-ray angiograms, of which $1,621$ are used for pre-training.
@@
LSF (Buades, Coll, and Morel 2005) [CVPR’05] $28.36$ $0.6300$
@@
DINO (Caron et al. 2021) [ICCV’21] $35.08$ $0.8840$
MAE (He et al. 2022) [CVPR’22] $35.11$ $0.8837$
DeblurringMIM (Kang et al. 2024) [MedIA’24] $35.10$ $0.8830$
HPM (Wang et al. 2025) [TPAMI’25] $35.10$ $0.8835$
VasoMIM [Ours] $35.14$ $0.8842$
Table S2: Comparison of state-of-the-art methods on CoronaryDominance. All methods are reimplemented using their official codebases. The best and second-best results are highlighted in red and yellow, respectively. Results are averaged over three random seeds.
Implementation Details
Vessel Segmentation. U-Net (Ronneberger, Fischer, and Brox 2015) is used as the segmentation decoder and fine-tuned for $500$ epochs on each dataset. All input images are resized to $224\times 224$. We employ the AdamW optimizer with an initial learning rate of $1e^{-4}$ and a weight decay of $0.05$. A cosine-annealing learning rate schedule with $T\_\max=500$ is applied.
More Results
@@
$80$ Frangi $27.06$ $\uparrow 15.00$ $25.10$ $\uparrow 0.64$
VasoMIM $68.30$ $\uparrow 0.33$ $84.29$ $\uparrow 0.23$
$92$ Frangi $41.30$ $\uparrow 29.24$ $64.01$ $\uparrow 39.55$
VasoMIM $\mathbf{68.85}$ $\mathbf{\uparrow 0.88}$ $\mathbf{84.49}$ $\mathbf{\uparrow 0.43}$
@@
Joint $68.78$ $\uparrow 0.48$ $84.06$ $\uparrow 0.10$
In Advance $\mathbf{68.85}$ $\mathbf{\uparrow 0.55}$ $\mathbf{84.49}$ $\mathbf{\uparrow 0.53}$
@@
$\mathcal{L}_{\rm w-rec.}$ $68.78$ $\uparrow 0.48$ $84.14$ $\uparrow 0.18$
$\mathcal{L}_{\rm cons.}$ $\mathbf{68.85}$ $\mathbf{\uparrow 0.55}$ $\mathbf{84.49}$ $\mathbf{\uparrow 0.53}$
@@
$\text{L}_{1}$ $68.81$ $\uparrow 0.51$ $84.32$ $\uparrow 0.36$
Dice $67.97$ $\downarrow 0.33$ $84.12$ $\uparrow 0.16$
Cross Entropy $\mathbf{68.85}$ $\mathbf{\uparrow 0.55}$ $\mathbf{84.49}$ $\mathbf{\uparrow 0.53}$
@@
Method # params (M) GFLOPs Speed (s/epoch)
MAE $111.65~(1.00\times)$ $9.43~(1.02\times)$ $54~(1.30\times)$
VasoMIM $111.91~(1.00\times)$ $9.59~(1.00\times)$ $70~(1.00\times)$

### v2

@@ Conclusion
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFC2415100, in part by the National Natural Science Foundation of China under Grant 62222316, Grant 62373351, Grant 82327801, Grant 62303463, in part by the Chinese Academy of Sciences Project for Young Scientists in Basic Research under Grant No. YSBR-104, in part by the Beijing Natural Science Foundation under Grant F252068, Grant 4254107, in part by Beijing Nova Program under Grant 20250484813, in part by China Postdoctoral Science Foundation under Grant 2024M763535, in part by the Postdoctoral Fellowship Program of CPSF under Grant GZC20251170.
@@ Acknowledgments
Pre-Training. (1) ARCADE (Popov et al. 2024): This dataset comprises two subsets for coronary vessel segmentation and atherosclerotic plaque localization. Each subset contains $1,500$ images, split into $1,000$ for training, $200$ for validation and $300$ for testing. Only the $1,000$ training images per subset are used for pre-training. (2) CADICA (Jiménez-Partinen et al. 2024): It comprises coronary X-ray angiography videos from $42$ patients, with frames of $512\times 512$ pixels and video lengths ranging from $1$ to $151$ frames. We select $6,594$ high-quality frames from these videos for pre-training. (3) Stenosis (Danilov et al. 2021): It offers X-ray angiography videos from $100$ patients with a total of $8,325$ frames. We use the $7,492$ training frames (from $90$ patients) for pre-training. (4) SYNTAX (Mahmoudi et al. 2025): This dataset includes $2,943$ X-ray angiograms from $231$ patients. All images are used for pre-training. (5) XCAD (Ma et al. 2021): This dataset provides $1,747$ X-ray angiograms, of which $1,621$ are used for pre-training.
@@
LSF (Buades, Coll, and Morel 2005) $28.36$ $0.6300$
@@
DINO (Caron et al. 2021) $35.08$ $0.8840$
MAE (He et al. 2022) $35.11$ $0.8837$
DeblurringMIM (Kang et al. 2024) $35.10$ $0.8830$
HPM (Wang et al. 2025) $35.10$ $0.8835$
VasoMIM $\bm{35.14}$ $\bm{0.8842}$
Table S2: Comparison of state-of-the-art methods on CoronaryDominance. The best results are highlighted in bold and the second-best results are underlined. Results are averaged over three random seeds.
@@
$80$ Frangi $27.06$ $25.10$
VasoMIM $68.30$ $84.29$
$92$ Frangi $41.30$ $64.01$
VasoMIM $\mathbf{68.85}$ $\mathbf{84.49}$
@@
Implementation Details
Vessel Segmentation. U-Net (Ronneberger et al. 2015) is used as the segmentation decoder and fine-tuned for $500$ epochs on each dataset. All input images are resized to $224\times 224$. We employ the AdamW optimizer with an initial learning rate of $1e^{-4}$ and a weight decay of $0.05$. A cosine-annealing learning rate schedule with $T\_\max=500$ is applied.
More Results
@@
Joint $68.78$ $84.06$
In Advance $\mathbf{68.85}$ $\mathbf{84.49}$
@@
$\mathcal{L}_{\rm w-rec.}$ $68.78$ $84.14$
$\mathcal{L}_{\rm cons.}$ $\mathbf{68.85}$ $\mathbf{84.49}$
@@
$\text{L}_{1}$ $68.81$ $84.32$
Dice $67.97$ $84.12$
Cross Entropy $\mathbf{68.85}$ $\mathbf{84.49}$
@@
Method # params (M) GFLOPs Speed (s/epoch)
MAE $111.65~(1.00\times)$ $18.86~(1.02\times)$ $54~(1.30\times)$
VasoMIM $111.91~(1.00\times)$ $19.17~(1.00\times)$ $70~(1.00\times)$